OpenStudy (anonymous):

The Ka of a monoprotic weak acid is 6.07 × 10-3. What is the percent ionization of a 0.179 M solution of this acid?

OpenStudy (vincent-lyon.fr):

Write the chemical equation of dissociation of your acid AH with water, with initial concentration \(c_o\), then consider that a fraction \(\alpha\) reacts with water. Concentration and Ka are big enough so that you can consider that \(1-\alpha\approx 1\) Write expression of Ka, then solve for \(\alpha\). You should find about 18.4%